One of the more sadistic rides at amusement parks is the “gravitron”. This is a vertical cylinder with a radius of 1.5 m. People stand inside the cylinder, against the wall. The cylinder spins, and once people reach a sufficient velocity the floor drops out from under them, leaving them stuck on the wall. If the minimum coefficient of static friction between the people and the wall is µs = 0.70, what must be the minimum linear velocity of the cylinder wall before the floor drops, if the people are to stick?
Answer #1

For circular motion, centripetal force is provided by the normal in this case while friction balances weight mg

N = mv^2/r

Friction = uN = mg

  umv^2/r = mg

0.70v^2 /1.5 = 9.8

v = sqrt(9.8*1.5/0.70)

= 4.583 m/s Answer